\begin{thebibliography}{ACRT18}

\bibitem[ACRT18]{ACRT18}
B.~Anthes, A.~Cattaneo, S.~Rollenske, and A.~Tomassini.
\newblock {$\partial\overline{\partial}$}-complex symplectic and {C}alabi-{Y}au
  manifolds: {A}lbanese map, deformations and period maps.
\newblock {\em Ann. Global Anal. Geom.}, 54(3):377--398, 2018.

\bibitem[BDIP02]{BDIP02}
J.~Bertin, J.-P. Demailly, L.~Illusie, and C.~Peters.
\newblock {\em Introduction to {H}odge theory}, volume~8 of {\em SMF/AMS Texts
  and Monographs}.
\newblock American Mathematical Society, Providence, RI; Soci\'{e}t\'{e}
  Math\'{e}matique de France, Paris, 2002.
\newblock Translated from the 1996 French original by James Lewis and Chris
  Peters.

\bibitem[Che22]{Chen22}
R.~Chen.
\newblock Moishezon deformations of manifolds in {F}ujiki class
  {$\mathcal{C}$}.
\newblock {\em J. Geom. Phys.}, 179:Paper No. 104603, 6, 2022.

\bibitem[CK99]{CK99}
David~A. Cox and Sheldon Katz.
\newblock {\em Mirror symmetry and algebraic geometry}, volume~68 of {\em
  Mathematical Surveys and Monographs}.
\newblock American Mathematical Society, Providence, RI, 1999.

\bibitem[Cle05]{Cle05}
H.~Clemens.
\newblock Geometry of formal {K}uranishi theory.
\newblock {\em Adv. Math.}, 198(1):311--365, 2005.

\bibitem[CMSP17]{CMP17}
J.~Carlson, S.~M\"{u}ller-Stach, and C.~Peters.
\newblock {\em Period mappings and period domains}, volume 168 of {\em
  Cambridge Studies in Advanced Mathematics}.
\newblock Cambridge University Press, Cambridge, 2017.
\newblock Second edition.

\bibitem[Fin24]{Fin24}
A.~Fino.
\newblock Canonical metrics in complex geometry.
\newblock {\em Boll. Unione Mat. Ital.}, 2024.
\newblock https://doi.org/10.1007/s40574-024-00421-y.

\bibitem[Fle81]{Fle81}
H.~Flenner.
\newblock Ein {K}riterium f\"{u}r die {O}ffenheit der {V}ersalit\"{a}t.
\newblock {\em Math. Z.}, 178(4):449--473, 1981.

\bibitem[FM06]{FM06}
D.~Fiorenza and M.~Manetti.
\newblock L-infinity algebras, {C}artan homotopies and period maps.
\newblock arXiv:math/0605297, 2006.

\bibitem[FM09]{FM09}
D.~Fiorenza and M.~Manetti.
\newblock A period map for generalized deformations.
\newblock {\em J. Noncommut. Geom.}, 3(4):579--597, 2009.

\bibitem[GGK13]{GGK13}
M.~Green, P.~Griffiths, and M.~Kerr.
\newblock {\em Hodge theory, complex geometry, and representation theory},
  volume 118 of {\em CBMS Regional Conference Series in Mathematics}.
\newblock Published for the Conference Board of the Mathematical Sciences,
  Washington, DC; by the American Mathematical Society, Providence, RI, 2013.

\bibitem[Gri68]{Gri68}
P.~Griffiths.
\newblock Periods of integrals on algebraic manifolds. ii: {Local} study of the
  period mapping.
\newblock {\em Amer. J. Math.}, 90(54):805--865, 1968.

\bibitem[HX24]{HX24}
Y.~Hu and W.~Xia.
\newblock Deformed {A}eppli cohomology: canonical deformations and jumping
  formulas.
\newblock preprint, 2024.

\bibitem[Kas21]{Kas21}
H.~Kasuya.
\newblock Introduction to variations of {H}odge structure over {S}asakian
  manifolds.
\newblock {\em Josai Mathematical Monographs}, 13:31--51, 2021.

\bibitem[Kir15]{Kir15}
T.~Kirschner.
\newblock {\em Period mappings with applications to symplectic complex spaces},
  volume 2140 of {\em Lecture Notes in Mathematics}.
\newblock Springer, Cham, 2015.

\bibitem[Kod86]{Kod86}
K.~Kodaira.
\newblock {\em Complex manifolds and deformation of complex structures}.
\newblock Springer-Verlag, 1986.

\bibitem[KP16]{KP16}
M.~Kerr and G.~Pearlstein, editors.
\newblock {\em Recent advances in {H}odge theory}, volume 427 of {\em London
  Mathematical Society Lecture Note Series}. Cambridge University Press,
  Cambridge, 2016.

\bibitem[KS60]{KS60}
K.~Kodaira and D.~C. Spencer.
\newblock On deformations of complex analytic structures. {III}. {S}tability
  theorems for complex structures.
\newblock {\em Ann. of Math. (2)}, 71:43--76, 1960.

\bibitem[LRY15]{LRY15}
K.~Liu, S.~Rao, and X.~Yang.
\newblock Quasi-isometry and deformations of {C}alabi-{Y}au manifolds.
\newblock {\em Invent. Math.}, 199(2):423--453, 2015.

\bibitem[LS18]{LS18}
K.~Liu and Y.~Shen.
\newblock Moduli spaces as ball quotients {I}, local theory.
\newblock arXiv:1810.10892, 2018.

\bibitem[LS20]{LS20}
K.~Liu and Y.~Shen.
\newblock Canonical sections of {Hodge} bundles on moduli spaces.
\newblock {\em J. Iranian Math. Soc.}, 1(1):97--115, 2020.

\bibitem[LSY09]{LSY09}
K.~Liu, X.~Sun, and S.-T. Yau.
\newblock Recent development on the geometry of the {Teichm\"uller} and moduli
  spaces of {Riemann} surfaces.
\newblock In {\em Geometry of Riemann surfaces and their moduli spaces}, volume
  XIV of {\em {Surveys in differential geometry}}, pages 221--259. 2009.

\bibitem[Man04]{Ma05}
M.~Manetti.
\newblock Lectures on deformations of complex manifolds (deformations from
  differential graded viewpoint).
\newblock {\em Rend. Mat. Appl. (7)}, 24(1):1--183, 2004.

\bibitem[Man22]{Man22}
M.~Manetti.
\newblock {\em Lie methods in deformation theory}.
\newblock Springer Monographs in Mathematics. Springer, Singapore, [2022]
  \copyright 2022.

\bibitem[Pop]{Popbook}
D.~Popovici.
\newblock {\em Non-{K\"{a}hler Hodge} Theory and Deformations of Complex
  Structures}.
\newblock available
  at~~{https://www.math.univ-toulouse.fr/~popovici/hodge-def.pdf}.

\bibitem[Pop19]{Pop19}
D.~Popovici.
\newblock Holomorphic deformations of balanced {Calabi-Yau}
  $\partial\bar{\partial}$-manifolds.
\newblock {\em Ann. Inst. Fourier (Grenoble)}, 69(2):673--728, 2019.

\bibitem[RZ18]{RZ18}
S.~Rao and Q.~Zhao.
\newblock Several special complex structures and their deformation properties.
\newblock {\em J. Geom. Anal.}, 28(4):2984--3047, 2018.

\bibitem[Sim91]{Sim91}
Carlos~T. Simpson.
\newblock The ubiquity of variations of {H}odge structure.
\newblock In {\em Complex geometry and {L}ie theory ({S}undance, {UT}, 1989)},
  volume~53 of {\em Proc. Sympos. Pure Math.}, pages 329--348. Amer. Math.
  Soc., Providence, RI, 1991.

\bibitem[Ste22]{Ste22}
J.~Stelzig.
\newblock Some remarks on the {S}chweitzer complex.
\newblock arXiv:2204.06027, 2022.

\bibitem[Tod89]{Tod89}
A.~N. Todorov.
\newblock The {Weil-Petersson} geometry of the moduli space of {SU($\geq3$)}
  {(Calabi-Yau)} manifolds i.
\newblock {\em Comm. Math. Phys.}, 126(2):325--346, 1989.

\bibitem[Voi02]{Voi02I}
C.~Voisin.
\newblock {\em Hodge theory and complex algebraic geometry. {I}}, volume~76 of
  {\em Cambridge Studies in Advanced Mathematics}.
\newblock Cambridge University Press, Cambridge, 2002.
\newblock Translated from the French original by Leila Schneps.

\bibitem[Voi05]{Voi05}
C.~Voisin.
\newblock Hodge theory and the topology of compact {K}\"ahler and complex
  projective manifolds.
\newblock In {\em Lecture Notes for the Seattle AMS Summer Institute}, 2005.

\bibitem[WX23]{WX23}
X.~Wan and W.~Xia.
\newblock Deformations of $(p,q)$-forms and degenerations of the
  {F}r\"{o}licher spectral sequence.
\newblock arXiv:2312.13692, 2023.

\bibitem[WZ24]{WZ20}
D.~Wei and S.~Zhu.
\newblock Two {A}pplications of the {$\partial \overline\partial$}-{H}odge
  {T}heory.
\newblock {\em Chinese Ann. Math. Ser. B}, 45(1):137--150, 2024.

\bibitem[Xia19]{Xia19deri}
W.~Xia.
\newblock Derivations on almost complex manifolds.
\newblock {\em Proc. Amer. Math. Soc.}, 147:559--566, 2019.
\newblock Errata in arXiv:1809.07443v3.

\bibitem[Xia21]{Xia19dBC}
W.~Xia.
\newblock On the deformed {B}ott-{C}hern cohomology.
\newblock {\em J. Geom. Phys.}, 166:Paper No. 104250, 19, 2021.

\bibitem[Xia22]{Xia19dDol}
W.~Xia.
\newblock Deformations of {Dolbeault} cohomology classes.
\newblock {\em Math. Z.}, 300:2931--2973, 2022.

\end{thebibliography}

\documentclass[11pt]{amsart}
\usepackage[mathscr]{euscript}
\usepackage{amssymb}

\usepackage{amscd}
\usepackage{amsmath, amssymb}
\usepackage{amsfonts}
\usepackage[colorlinks,linkcolor=blue,citecolor=blue, pdfstartview=FitH]{hyperref}
\usepackage[all]{xy}

\numberwithin{equation}{section}

\theoremstyle{plain}
\newtheorem{thm}{Theorem}[section]
\newtheorem{theorem}[thm]{Theorem}

\newtheorem{corollary}[thm]{Corollary}
\newtheorem{proposition}[thm]{Proposition}

\theoremstyle{definition}

\newtheorem{definition}[thm]{Definition}

\newtheorem{defn-thm}[thm]{Definition-Theorem}

\newcommand{\im}{{ \mathrm{im}\,}}

\newcommand{\C}{{ \mathbb{C} }}
\newcommand{\Z}{{ \mathbb{Z} }}

\newcommand{\G}{{\textrm{Grass}\,}}
\newcommand{\p}{{ \partial }}
\newcommand{\pb}{{ \bar{\partial} }}

\newcommand{\wg}{{\,\wedge\,}}
\newcommand{\Hom}{{ \textrm{Hom}\,}}

\begin{document}

\title{Variation of Hodge structures for non-K\"ahler manifolds}
\author{Wei Xia}
\address{Wei Xia, Mathematical Science Research Center, Chongqing University of Technology, Chongqing, P.R.China, 400054.} \email{xiawei@cqut.edu.cn, xiaweiwei3@126.com}
\subjclass[2020]{32G05, 32G20, 55T05}
 \keywords{variation of Hodge structures, non-K\"ahler manifolds, period maps.}
 
\thanks{This work was supported by the National Natural Science Foundation of China No. 11901590 and Natural Science Foundation of Chongqing (China) No. CSTB2022NSCQ-MSX0876 and Chongqing Natural Science Foundation Innovation Development Joint Fund No. CSTB2024NSCQ-LZX0040.}
\date{\today}

\begin{abstract}
In this note, we discuss unpolarized, complex variation of Hodge structures for non-K\"ahler manifolds. In particular, given a holomorphic family of compact complex manifolds whose central fiber satisfies: the inclusions $F^{p}A^{p+q+1}(X)\hookrightarrow A^{p+q+1}(X), F^{p}A^{p+q}(X)\hookrightarrow A^{p+q}(X)$ are injective in cohomology, it is shown that the period map is holomorphic and transversal.
\end{abstract}

\maketitle

%\tableofcontents
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Introduction}
The theory of variation of Hodge structures (VHS for short) is an extensively studied object in complex geometry and related subjects (see \cite{GGK13,CMP17,KP16,CK99} and the references therein). The purpose of this note is to extend some parts of Griffiths' classical theory of VHS \cite{Gri68} to the non-K\"ahler setting, see \cite{Sim91,Kir15,Kas21} for extensions in other directions. There are two motivations to do this, the first is the works of Popovici and Anthes-Cattaneo-Rollenske-Tomassini \cite{Pop19,ACRT18}, where they showed local Torelli theorem holds for Clabi-Yau $\p\pb$-manifolds and $\p\pb$-complex symplectic manifolds, respectively; the second is that we want to understand Voisin's proof (see \cite{Voi05}) on a density criterion for complex projective manifolds\footnote{This has been generalized by Rutong Chen \cite{Chen22} to Moishezon manifolds by using the results obtained in this note.}.

Let $X$ be a compact complex manifold, there is the \emph{Hodge filtration on forms}
\[
F^pA^{p+q}(X)=\bigoplus_{i=0}^{\infty}A^{p+i,q-i}(X),\quad p,q\in\Z,
\]
where $A^{p,q}(X)$ is the space of smooth complex valued differential forms on $X$. Following \cite{WX23}, elements in $F^pA^{p+q}(X)$ will be called \textit{filtered $(p,q)$-forms on $X$}. Since Hodge filtrations are compatible with the exterior differential $d$, it induces \emph{Hodge filtrations on the de Rham cohomology}:
\[
F^pH^{p+q}(X):=\im \left( \frac{\ker d\cap F^pA^{p+q}(X) }{\im d\cap F^pA^{p+q}(X) }\longrightarrow H^{p+q}(X,\C) \right),\quad p,q\in\Z.
\]
As in the classical theory, we are interested in the variations of $F^pH^{p+q}(X)$ when the complex structure on $X$ deforms. For this purpose, we first establish a deformation theory for filtered $(p,q)$-forms.

More precisely, let
\[
\pi: (\mathcal{X}, X)\to (B,0)
\]
be a complex analytic family over a small polydisc $B\subset\C^m$ such that for each $t\in B$ the complex structure on $X_t$ is represented by a Beltrami differential $\phi(t)$. Given
\[
\alpha_0\in\ker d\cap F^{p}A^{p+q}(X),
\]
and $T\subseteq B$, which is an analytic subset of $B$ containing $0$, a \textit{deformation} of $\alpha_0$ (w.r.t. $\pi$ ) on $T$ is a family of forms
\[
\alpha (t)\in F^{p}A^{p+q}(X),
\]
such that
\begin{itemize}
  \item[(1)] $\alpha (t)$ is holomorphic in $t$ and $\alpha(0) = \alpha_0$;
  \item[(2)] for any $t\in T$, we have
  \begin{equation}\label{eq-ext-eq}
  d_{\phi(t)}\alpha (t)=0.
  \end{equation}
\end{itemize}
We say \textit{the deformation of filtered $(p,q)$-forms on $X$ are unobstructed} (w.r.t. $\pi$) if for any $\alpha_0\in\ker d\cap F^{p}A^{p+q}(X)$ there is a deformation of $\alpha_0$ on $B$.

For any given closed filtered $(p,q)$-form $\alpha_0\in\ker d\cap F^{p}A^{p+q}(X)$, the deformations $\alpha(t)$ of $\alpha_0$ is by no means unique. Motivated by the deformation theory of Dolbeault cohomology classes \cite{Xia19dDol}, we are led to consider canonical deformations, i.e. we will construct deformations $\alpha(t)$ by using Hodge decomposition for filtered forms with respect to a fixed Hermitian metric on $X$. It may happen that $\alpha_0$ is not $d$-exact while $\alpha(t)$ is $d_{\phi(t)}$-exact for some $t$, i.e. $\alpha(t)$ is not a faithful deformation in the sense of Definition \ref{def-deformation-Dol-Zrpq}. As in the case of deformations of Dolbeault cohomology, this together with possible obstructions are exactly the causes which make $\dim F^pH^{p+q}(X_t)$ jumps at $t=0$. In Theorem \ref{thm-jumping-formula}, we will prove the following: for any $0\leq p\leq k$ and $t\in B$, set
\[
v^{p,k}_t:=\dim \ker\triangle_{p}\cap F^{p}A^{k}(X)-\dim F^pA^{k}(X)\cap\ker d_{\phi(t)}\cap\ker d^*_p \geq 0,
\]
where $d^*_p=d^*-\p^*\Pi^{p,\bullet-p}$ is the formal adjoint of $d$ on $F^{p}A^{k}(X)$ (c.f. Subsection \ref{subsec-Hodge-decomp}), then we have $(v^{p,-1}_t:=0)$
\begin{equation}\label{eq-jumping-formula-0}
\dim F^pH^{k}(X)=\dim F^pH^{k}(X_t)+v^{p,k}_t+v^{p,k-1}_t.
\end{equation}
This may be called \emph{the jumping formula for filtered de Rham cohomology}, see \cite{Xia19dDol,HX24} for versions of Dolbeault/Bott-Chern/Aeppli cohomology. These formula may be viewed as a refinement of the classical result obtained by Kodaira-Spencer (see formula $(43)$ in \cite[pp.\,68]{KS60}). In their formula, $v^q_t$ is defined via the spectrum of Laplacian operators. Here, the integers $v^{p,k}_t$ have significant geometric meanings, roughly speaking, it measures the size of classes in $F^pH^{k}(X)$ whose canonical deformation does not exist on $t$. As direct applications of the jumping formula \eqref{eq-jumping-formula-0}, we get the following
\begin{corollary}
The following holds:
\begin{enumerate}
  \item $\dim F^pH^{k}(X_t)$ is independent of $t\in B$ if and only if the deformations of filtered $(p,k-p)$-forms and filtered $(p,k-1-p)$-forms are canonically unobstructed;
  \item For any $0\leq p\leq k$, the alternating sum $\sum_{i=0}^k \dim (-1)^{k-i}F^pH^{i}(X_t)$ and $\dim E^{0,q}_1(X_t)/E^{0,q}_\infty(X_t)$ are upper semi-continuous function of $t\in B$ (in analytic Zariski topology). Moreover, the deformations of filtered $(p,k-p)$-forms are canonically unobstructed if and only if $\sum_{i=0}^k \dim (-1)^{k-i}F^pH^{i}(X_t)$ is independent of t;
  \item The notion of canonically unobstructedness for filtered forms are independent of the choices of Hermitian metrics.
\end{enumerate}
\end{corollary}
In $(2)$, $E^{p,q}_r(X_t)$ are the spaces in the Fr\"olicher spectral sequence of $X_t$. The upper semi-continuity of $\dim E^{0,q}_1(X_t)/E^{0,q}_\infty(X_t)$ for $q=1$ or $n-1$ was shown by Stelzig \cite{Ste22} (in Euclidean topology). On the other hand, it was shown by Flenner \cite{Fle81} that for any $q$, $\sum_{i=0}^{q}(-1)^{q-i}\dim H^{i}(X_t,\mathcal{E}\mid_{X_t})$ is upper semi-continuous in analytic Zariski topology, where $\mathcal{E}$ is a coherent analytic sheaf on the total space $\mathcal{X}$, see also \cite[Thm.\,5.10]{BDIP02}. Notice also that when $p=0, k=2\dim_\C X$, the alternating sum $\sum_{i=0}^k \dim (-1)^{k-i}F^pH^{i}(X_t)=\sum_{i=0}^k \dim (-1)^{i}H^{i}(X_t)$ is just the topological Euler number of $X$.

Let $\pi: (\mathcal{X}, X)\to (B,0)$ be a complex analytic family over a small polydisc $B\subset\C^m$ such that for each $t\in B$ the complex structure on $X_t$ is represented by a Beltrami differential $\phi(t)$. Following \cite[Chp.\,10]{Voi02I}, the \emph{period map} is defined as follows
\begin{equation}\label{eq-period-map-0}
\Phi^{p,k}:B\longrightarrow \G (f^{p,k}, H^{k}(X,\C)),\quad t\longmapsto F^pH^{k}(X_t),
\end{equation}
where $f^{p,k}:=\dim F^pH^{k}(X)$. Our main result is the following:
\begin{theorem}[=Theorem \ref{thm-period-map}]\label{thm-period-map-0}
Suppose $X$ is equipped with a Hermitian metric. Assume
\begin{equation}\label{eq-assumption-1}
  F^{p}A^{p+q+1}(X)\cap dA^{p+q}(X) =dF^{p}A^{p+q}(X),
\end{equation}
and
\begin{equation}\label{eq-assumption-2}
  F^{p}A^{p+q}(X)\cap dA^{p+q-1}(X) =dF^{p}A^{p+q-1}(X).
\end{equation}
The following holds:
\begin{enumerate}
  \item The period map $\Phi^{p,k}$ is holomorphic;
  \item Griffiths transversality: the tangent map
  \[
  d\Phi^{p,k}_0:T_0B\longrightarrow  \Hom(F^pH^{k}(X), H^{k}(X,\C)/F^pH^{k}(X))
  \]
  has values in $\Hom(F^pH^{k}(X), F^{p-1}H^{k}(X)/F^pH^{k}(X))$;
\end{enumerate}
If furthermore $\mathcal{H}_p=\mathcal{H}_{p-1}\mid_{F^{p}A^{\bullet}(X)}$ and $\dim B=1$, then the tangent map of $\Phi^{p,k}$ is given by
\[
d\Phi^{p,k}_0(\frac{\p}{\p t}\Big|_{t=0})([\alpha_0])=[\mathcal{H}_{p-1}i_{\phi_1}\alpha_0]\in F^{p-1}H^{k}(X),\quad [\alpha_0]\in F^pH^{k}(X),
\]
where $t$ is a local coordinate function on $B$ and $\phi_1$ is the first order coefficient in $\phi(t)=\sum_{j\geq 1}\phi_jt^j$.
\end{theorem}
The assumption \eqref{eq-assumption-1} says the inclusion $F^{p}A^{p+q+1}(X)\hookrightarrow A^{p+q+1}(X)$ is injective in cohomology. It holds for all $(p,q)$ if and only if the Fr\"olicher spectral sequence of $X$ degenerate at the first page (i.e. $E_1\cong E_\infty$), see Corollary C.6.7. in \cite{Man22}. Furthermore, in terms of the differentials $d_r^{p,q}:E_r^{p,q}(X)\longrightarrow E_r^{p+r,q-r+1}(X)$, \eqref{eq-assumption-1} is equivalent to $\bigoplus_{r\geq i\geq 1}d_r^{p-i,q+i}=0$, see Corollary B.6. in \cite{WX23}.

A main ingredient to prove this theorem is the fact that the exponential operator preserves Hodge filtrations, see Theorem \ref{thm-exp-FpH}. The period map defined by \eqref{eq-period-map-0} concerns unpolarized, complex VHS for non-K\"ahler manifolds. It is an interesting problem to extend $\Phi^{p,k}$ to the polarized version. Unfortunately, in our case as K\"ahler metrics are not assumed to exist, the Hodge-Riemann bilinear relations and hard Lefschetz theorem are not available. Moreover, it is perhaps worthwhile to point out that if the Hermitian metric $h$ is K\"ahler, the condition $\mathcal{H}_p=\mathcal{H}_{p-1}\mid_{F^{p}A^{\bullet}(X)}$ automatically holds. It seems an interesting question to determine the type of special Hermitian metrics (see \cite{Fin24,Popbook} and the references therein for detailed discussions) that make $\mathcal{H}_p=\mathcal{H}_{p-1}\mid_{F^{p}A^{\bullet}(X)}$ hold.
\vskip 1\baselineskip \textbf{Acknowledgements.} We would like to thank Prof. Kefeng Liu for critical comments. Many thanks to Xueyuan Wan, Rutong Chen, Yang Shen and Ya Deng for stimulating discussions.
\section{The exponential operator} \label{}
The exponential operator\footnote{An implicit form of the exponential operator was used by Griffiths in \cite[Prop.\,1.11]{Gri68}.}
\[
e^{i_{\phi(t)}}:=\sum_{k=0}^{\infty} \frac{i_{\phi(t)}^k}{k!}
\]
which was introduced in the work of Todorov (see Remark $2.4.9.$ in ~\cite[pp.\,339]{Tod89}) turns out to be very useful for the studying of deformations of complex manifolds, see e.g. \cite{Cle05,LSY09,LRY15,RZ18,Xia19dBC,LS20}.
Let $\pi: (\mathcal{X}, X)\to (B,0)$ be a complex analytic family in the sense of Kodaira-Spencer, that is, $\pi$ is a holomorphic submersion which is surjective and proper over the connected complex manifold $B$. Assume $B$ is a small polydisc (centered in $0$) with coordinates $t=(t_1,t_2,\cdots)$, then by Ehresmann's theorem we have the following commutative diagram
\[
  \xymatrix{
  X\times B \ar[dr] \ar[r]^-{F}
                & \mathcal{X} \ar[d]^-{\pi}  \\
                & B~, }
\]
where $F$ is a diffeomorphism. For any $t\in B$ set $f_t:=F\mid_{X\times \{t\}}$, then $f_t$ is a diffeomorphism from $X$ to $X_t$. We use the isomorphism
\[
f_t^*:A^\bullet(X_t)\longrightarrow A^\bullet(X)
\]
to identify forms/cohomologies on $X_t$ with forms/cohomologies on $X$.
Let $z^1,\cdots, z^n$ and $w^1,\cdots, w^n$ be holomorphic coordinates on $X$ and $X_t$ respectively, the Beltrami differential can be defined by
\begin{equation}\label{Beltrami differential coord}
\phi(t):=\left(\frac{\partial w}{\partial z}\right)^{-1}_{\alpha\gamma}\frac{\partial w^{\gamma}}{\partial \bar{z}^{\beta}} d\bar{z}^{\beta}\otimes\frac{\partial}{\partial z^{\alpha}}\in A^{0,1}(X, T^{1,0})~,
\end{equation}
where by abuse of notations, we write $w^i=f_t^*w^i=w^i\circ f_t$ for each $i=1,2,\cdots,n$.
From
\begin{equation}\label{dw in terms of dz}
d w^{\beta}=\frac{\partial w^{\beta}}{\partial z^{\alpha}}dz^{\alpha}+\frac{\partial w^{\beta}}{\partial \bar{z}^{\alpha}}d\bar{z}^{\alpha}
=\frac{\partial w^{\beta}}{\partial z^{\alpha}}(1+i_{\phi(t)})dz^{\alpha},
\end{equation}
and
\[
e^{i_{\phi(t)}} (dz^{i_1}\wg\cdots\wg dz^{i_p})=(dz^{i_1}+i_{\phi(t)}dz^{i_1}) \wg \cdots \wg (dz^{i_p}+i_{\phi(t)}dz^{i_p}),
\]
we see that
\begin{equation}\label{eq-exp}
e^{i_{\phi(t)}}~: A^{p,0}(X)\longrightarrow A^{p,0}(X_t).
\end{equation}
We will use the following formula (see e.g. \cite{LRY15,FM06,Xia19deri} and \cite[pp.\,78]{Ma05}):
\begin{equation}\label{eq-LRY15}
d_{\phi(t)}:=e^{-i_{\phi(t)}}de^{i_{\phi(t)}} = \partial + \bar{\partial}_{\phi(t)}=\partial + \bar{\partial}-\mathcal{L}_{\phi(t)}^{1,0}~,
\end{equation}
where $\mathcal{L}_{\phi(t)}^{1,0}:=i_{\phi(t)}\p-\p i_{\phi(t)}$ is the Lie derivative.
A remarkable property about the exponential operator (as observed in \cite{FM06,FM09,WZ20}) is that while it does not preserve all the types of $(p,q)$-forms, it does map elements in a filtration on $X$ to elements in the corresponding filtration on $X_t$:
\begin{equation}\label{eq-weidingchang}
e^{i_{\phi(t)}}~: F^pA^{k}(X)=\oplus_{p\leq\lambda\leq n} A^{\lambda,k-\lambda}(X)\longrightarrow F^pA^{k}(X_t),
\end{equation}
where $n=\dim_\C X$. In fact, let $\varphi\in A^{\lambda,k-\lambda}(X)$ with $p\leq\lambda\leq k$, locally we write $\varphi=\varphi_{IJ}dz^I\wedge d\bar{z}^J$ where $dz^I = dz^{i_1}\wedge\cdots\wedge dz^{i_\lambda}$ and $d\bar{z}^J = d\bar{z}^{j_1} \wedge\cdots\wedge d\bar{z}^{j_{k-\lambda}}$, then
\begin{align*}
&e^{i_{\phi(t)}}\varphi\\
=&\varphi_{IJ}(e^{i_{\phi(t)}}dz^I)\wedge d\bar{z}^J\\
=&\varphi_{IJ}(e^{i_{\phi(t)}}dz^I)\wg (\frac{\p\bar{z}^{j_1}}{\p w^l}dw^l+\frac{\p\bar{z}^{j_1}}{\p \bar{w}^l}d\bar{w}^l)\wg\cdots\wg (\frac{\p\bar{z}^{j_{k-\lambda}}}{\p w^l}dw^l+\frac{\p\bar{z}^{j_{k-\lambda}}}{\p \bar{w}^l}d\bar{w}^l),
\end{align*}
which is clearly an element in $F^pA^{k}(X_t)$. It follows that \cite{WX23}
\begin{equation}\label{eq-wdc-dc}
e^{i_{\phi(t)}}:(F^\bullet A^{\bullet}(X),d_{\phi(t)} )\to (F^\bullet A^{\bullet}(X_t),d)
\end{equation}
is an isomorphism of filtered complexes.

\section{Deformations of filtered forms} \label{Deformations of filtered froms}
In this section, we establish a deformation theory for filtered forms which is modeled on \cite{Xia19dDol}.
\subsection{Hodge decomposition for filtered forms}\label{subsec-Hodge-decomp}
Assume $X$ has been equipped with a fixed Hermitian metric. There is a Hodge decomposition for filtered forms (c.f. Appendix A of \cite{WX23}). In fact, set
\[
d^*_p=\Pi^{\geq p}d^*:F^{p}A^{\bullet}(X)\longrightarrow F^{p}A^{\bullet}(X),
\]
where $\Pi^{\geq p}:A^{\bullet}(X)\to F^pA^{\bullet}(X)$ is the linear projection onto $F^pA^{\bullet}(X)$. It turns out $d^*_p$ is the formal adjoint of $d$ for filtered forms, namely,
\[
(d\alpha,\beta)=(\alpha,d^*_p\beta),\quad \text{for~any}~\alpha\in F^pA^{\bullet}(X), \beta\in F^pA^{\bullet+1}(X).
\]
Notice that $d^*_p=d^*-\p^*\Pi^{p,\bullet-p}$ where $\Pi^{p,q}:A^{\bullet}(X)\to A^{p,q}(X)$ is the linear projection onto $A^{p,q}(X)$. So we have $d^*_p\mid_{F^{p+1}A^{\bullet}(X)}=d^*$.
The \emph{Laplacian operator for filtered forms} is defined by
\[
\triangle_p=dd^*_p+d^*_pd:F^{p}A^{\bullet}(X)\longrightarrow F^{p}A^{\bullet}(X),
\]
then there is the \emph{Green operator} of $\triangle_p$, such that
\[
1=\mathcal{H}_p+\triangle_pG_p=\mathcal{H}_p+G_p\triangle_p,\quad\text{on}~F^pA^{\bullet}(X),
\]
where $\mathcal{H}_p$ is the projection onto the harmonic space $\ker\triangle_{p}\cap F^{p}A^{k}(X)$.
Now if $\pi: (\mathcal{X}, X)\to (B,0)$ be a complex analytic family over a small polydisc $B\subset\C^m$ such that for each $t\in B$ the complex structure on $X_t$ is represented by a Beltrami differential $\phi(t)$. We have the following:
\begin{itemize}
  \item For any $t\in B$, set
  \[
  \triangle_{p\phi(t)}=d_{p\phi(t)}^*d_{\phi(t)}+d_{p\phi(t)} d_{p\phi(t)}^*:F^{p}A^{\bullet}(X)\longrightarrow F^{p}A^{\bullet}(X),
  \]
  where $d_{p\phi(t)}^*=\Pi^{\geq p}d_{\phi(t)}^*$ and $d_{\phi(t)}^*$ is the formal adjoint of $d_{\phi(t)}$;
  \item There is a Green operator $G_{p\phi(t)}:F^{p}A^{\bullet}(X)\longrightarrow F^{p}A^{\bullet}(X)$ such that
  \[
  1=\mathcal{H}_{p\phi(t)}+\triangle_{p\phi(t)}G_{p\phi(t)}
   =\mathcal{H}_{p\phi(t)}+G_{p\phi(t)}\triangle_{p\phi(t)},
  \]
  where $\mathcal{H}_{p\phi(t)}$ is the harmonic projection operator.
\end{itemize}
It is clear that $\triangle_{p\phi(t)}$ is a differentiable family of formally self-adjoint, elliptic differential operators in the sense of Kodaira-Spencer \cite{KS60} (see also \cite{Kod86}). As a result, by Kodaira-Spencer's upper semi-continuity theorem, we get that for any $t\in B$,
\[
\dim F^pH^{k}(X)=\dim \ker \triangle_{p}\geq \dim \ker \triangle_{p\phi(t)}=\dim F^pH^{k}_{d_{\phi(t)}}(X)= \dim F^pH^{k}(X_t),
\]
where $F^pH^{k}_{d_{\phi(t)}}(X):=\frac{\ker d_{\phi(t)}\cap F^pA^{k}(X)}{\im d_{\phi(t)}\cap F^pA^{k}(X)}$ and it follows from \eqref{eq-weidingchang} that
\[
F^pH^{k}_{d_{\phi(t)}}(X)\cong F^pH^{k}(X_t).
\]
As a result, we have the following:
\begin{proposition}\label{prop-upper-semicontinuous}
Let $\pi: (\mathcal{X}, X)\to (B,0)$ be a complex analytic family over a small polydisc $B\subset\C^m$. Then $\dim F^pH^{\bullet}(X_t)$ is an upper semi-continuous function of $t\in B$.
\end{proposition}
An immediate consequence is that for any $q\geq0$, $\dim E^{0,q}_\infty$ is lower semi-continuous in deformation of complex structures. In fact, from
\begin{equation}\label{eq-E0q}
E^{0,q}_\infty(X)\cong \frac{F^0H^{q}(A^{\bullet}(X),d)}{F^{1}H^{q}(A^{\bullet}(X),d)}\cong \frac{H^{q}_{dR}(X)}{F^{1}H^{q}(X) },
\end{equation}
we get $\dim E^{0,q}_\infty(X)=b^q-\dim F^{1}H^{q}(X)$, where $b^q=\dim H^{q}_{dR}(X)$ is the $q$-th Betti number of $X$. By using Proposition \ref{prop-upper-semicontinuous}, we get the following
\begin{corollary}
$\dim E^{0,q}_\infty$ is lower semi-continuous in deformation of complex structures. $\dim E^{0,q}_1/E^{0,q}_\infty$ is upper semi-continuous in deformation of complex structures.
\end{corollary}
\begin{proof}The second statement follows since $\dim E^{0,q}_1=h^{0,q}$, the $(0,q)$-th Hodge number, is upper semi-continuous and the sum of two upper semi-continuous function is still upper semi-continuous.
\end{proof}
There is a strengthened version of this result which holds in analytic Zariski topology, see Corollary \ref{Coro-123}.
\subsection{Construction of deformations}
We write $\phi(t)=\sum_{i\geq1}\phi_i$.
\begin{proposition}\label{prop-unique-power-seris-solution}
\begin{enumerate}
  \item For any $t\in B$, the following natural homomorphism induced by inclusion $\ker d_{\phi(t)}\cap\ker d^*_p\subset\ker d_{\phi(t)}$ is an isomorphism:
\[
\frac{F^pA^{k}(X)\cap\ker d_{\phi(t)}\cap\ker d^*_p}{F^pA^{k}(X)\cap\im d_{\phi(t)}\cap\ker d^*_p}\longrightarrow
\frac{F^pA^{k}(X)\cap\ker d_{\phi(t)}}{F^pA^{k}(X)\cap\im d_{\phi(t)}}=F^pH^{k}_{d_{\phi(t)}}(X).
\]
  \item For any given $\alpha\in F^{p}A^{k}(X)$, if $d_{\phi(t)}\alpha = d\alpha - \mathcal{L}_{\phi(t)}^{1,0}\alpha =0$ and $d^*_p\alpha=0$, then we must have
\[
\alpha = \mathcal{H}_{p}\alpha + d^*_p G_p\mathcal{L}_{\phi(t)}^{1,0}\alpha .
\]
  \item For any fixed $\alpha_0\in\ker\triangle_{p}\cap F^{p}A^{k}(X)$, the equation
\begin{equation}\label{Kuranishi eq}
\alpha=\alpha_0 + d^*_p G_p\mathcal{L}_{\phi(t)}^{1,0}\alpha,
\end{equation}
has an unique solution given by $\alpha=\alpha(t)=\sum_{k\geq0} \alpha_k$ and
\[
\alpha_k=d^*_p G_p\sum_{i+j=k} \mathcal{L}_{\phi_i}^{1,0}\alpha_{j}\in F^{p}A^{p+q}(X),\quad k>0,
\]
which converges for $|t|$ small.
\item Let $\alpha$ be a solution of the equation \eqref{Kuranishi eq}.
Then for any $t\in B$, we have
\begin{equation}\label{eq-equivalence-harmonic part vanish}
d_{\phi(t)}\alpha=0 \Leftrightarrow \mathcal{H}_p\mathcal{L}_{\phi(t)}^{1,0}\alpha =0.
\end{equation}
\item Set
\begin{align*}
\hat{g}_t:\ker\triangle_{p}\cap F^{p}A^{k}(X)&\longrightarrow \ker d^*\cap\im d_{\phi(t)}\cap F^{p}A^{k+1}(X)\\
x_0&\longmapsto d_{\phi(t)}x(t),
\end{align*}
where $x(t)$ is the unique solution of $x(t)=x_0+d^*_p G_p\mathcal{L}_{\phi(t)}^{1,0}x(t)$. Then the following holds:
\[
\dim\ker\triangle_{p}\cap F^{p}A^{k}(X)=\dim\left(\ker d_p^*\cap\ker d_{\phi(t)}\right)^k+
\dim \left(\ker d_p^*\cap\im d_{\phi(t)}\right)^{k+1},
\]
where
\begin{itemize}
  \item $\left(\ker d_p^*\cap\ker d_{\phi(t)}\right)^k:=\ker d_p^*\cap\ker d_{\phi(t)}\cap F^{p}A^{k}(X)$;
  \item $\left(\ker d_p^*\cap\im d_{\phi(t)}\right)^{k+1}:=\left(\ker d_p^*\cap\im d_{\phi(t)}\right)^{k+1}\cap F^{p}A^{k+1}(X)$.
\end{itemize}
\end{enumerate}
\end{proposition}
\begin{proof}
This follows from minor modifications of the arguments in \cite[Sec.\,4]{Xia19dDol}.
\end{proof}

\begin{definition}\label{def-V_t^k}
For any $t\in B$ and a vector subspace $V\subseteq \ker\triangle_{p}\cap F^{p}A^{k}(X)$, we set
\begin{align*}
V_{t}^{p,k}:=
\Big\{ \alpha_0\in V \mid~~ &d_{\phi(t)}\alpha(t)=0~\text{where}
~\alpha(t)\\
&\text{is the unique solution of}~\alpha (t)=\alpha_0+d^*_p G_p\mathcal{L}_{\phi(t)}^{1,0}\alpha(t)\Big\}.
\end{align*}
The set $\{t\in B\mid\dim V_{t}^{p,k}\geq n\}$ is an analytic subset of $B$ for any $n\in \mathbb{N}$.
\end{definition}

\begin{definition}\label{def-f_t-g_t}
For any $t\in B$, we set
\begin{align*}
f_t: &V_{t}^{p,k} \longrightarrow \frac{F^pA^{k}(X)\cap\ker d_{\phi(t)}\cap\ker d^*_p}{F^pA^{k}(X)\cap\im d_{\phi(t)}\cap\ker d^*_p}\cong F^pH^{k}_{d_{\phi(t)}}(X),\\
&\alpha_0\longmapsto \alpha(t)=\sum_{k\geq0} \alpha_k,~\text{where}~\alpha_{k}=d^*_p G_p\sum_{i+j=k}\mathcal{L}_{\phi_i}^{1,0}\alpha_{j},~\text{for~any}~ k> 0.
\end{align*}
\end{definition}

\begin{proposition}\label{prop V-E-t}
If $V=\ker\triangle_{p}\cap F^{p}A^{k}(X)$, then $f_t$ is surjective with
\[
\ker f_t\cong F^pA^{k}(X)\cap\im d_{\phi(t)}\cap\ker d^*_p.
\]
\end{proposition}
\begin{proof}By Proposition \ref{prop-unique-power-seris-solution}, the following map
\begin{align*}
\tilde{f}_t: &V_{t}^{p,k} \longrightarrow F^pA^{k}(X)\cap\ker d_{\phi(t)}\cap\ker d^*_p,\\
&\alpha_0\longmapsto \alpha(t)=\sum_{k\geq0} \alpha_k,~\text{where}~\alpha_{k}=d^*_p G_p\sum_{i+j=k}\mathcal{L}_{\phi_i}^{1,0}\alpha_{j},~\text{for~any}~ k\neq 0,
\end{align*}
is an isomorphism.
\end{proof}

\begin{theorem}\label{thm-jumping-formula}
For any $0\leq p\leq k$ and $t\in B$, set
\[
v^{p,k}_t:=\dim \ker\triangle_{p}\cap F^{p}A^{k}(X)-\dim F^pA^{k}(X)\cap\ker d_{\phi(t)}\cap\ker d^*_p \geq 0,
\]
then we have $(v^{p,-1}_t:=0)$
\begin{equation}\label{eq-jumping-formula}
\dim F^pH^{k}(X)=\dim F^pH^{k}(X_t)+v^{p,k}_t+v^{p,k-1}_t.
\end{equation}
\end{theorem}
\begin{proof}First, by Proposition \ref{prop V-E-t}, we have
\[
\dim V_{t}^{p,k}-\dim F^pA^{k}(X)\cap\im d_{\phi(t)}\cap\ker d^*_p=\dim F^pH^{k}_{d_{\phi(t)}}(X)=\dim F^pH^{k}(X_t).
\]
Combining this with $(5)$ of Proposition \ref{prop-unique-power-seris-solution}, we have
\[
\dim V_{t}^{p,k}-\dim\ker\triangle_{p}\cap F^{p}A^{k-1}(X)+\dim\left(\ker d_p^*\cap\ker d_{\phi(t)}\right)^{k-1}=\dim F^pH^{k}(X_t),
\]
which by the definition of $v^{p,k-1}_t$ implies
\[
\dim V_{t}^{p,k}=\dim F^pH^{k}(X_t)+v^{p,k-1}_t.
\]
On the other hand, we know that
\begin{equation}\label{eq-Vt-vt}
\dim V_{t}^{p,k}=F^pA^{k}(X)\cap\ker d_{\phi(t)}\cap\ker d^*_p=\dim \ker\triangle_{p}\cap F^{p}A^{k}(X)-v^{p,k}_t.
\end{equation}
Hence, \eqref{eq-jumping-formula} follows.
\end{proof}
Forms in $F^pA^{p+q}(X)=\bigoplus_{i=0}^{\infty}A^{p+i,q-i}(X)$ will be called \textit{filtered $(p,q)$-forms on $X$}.
\begin{definition}\label{def-deformation-Dol-Zrpq}
Given
\[
\alpha_0\in\ker d\cap F^{p}A^{p+q}(X),
\]
and $T\subseteq B$, which is an analytic subset of $B$ containing $0$, a \textit{deformation} of $\alpha_0$ (w.r.t. $\pi$ ) on $T$ is a family of forms
\[
\alpha (t)=\alpha^{p,q}(t)+\alpha^{p+1,q-1}(t)+\cdots+\alpha^{n,p+q-n}(t)
\in F^{p}A^{p+q}(X),
\]
such that
\begin{itemize}
  \item[(1)] $\alpha (t)$ is holomorphic in $t$ and $\alpha(0) = \alpha_0$;
  \item[(2)] for any $t\in T$, we have
  \begin{equation}\label{eq-ext-eq}
  d_{\phi(t)}\alpha (t)=0.
  \end{equation}
\end{itemize}
By a \emph{canonical deformation of $\alpha_0$} w.r.t. $\pi$, we mean a deformation of the form
\begin{equation}\label{eq-canonical}
\alpha(t)=\sum_{k\geq0} \hat{\alpha}_k+d_{\phi(t)}\beta_0,
\end{equation}
where $\beta_0\in F^{p}A^{p+q-1}(X)$ satisfies $d\beta_0=\alpha_0-\mathcal{H}_p\alpha_0$ and $\sum_{k\geq0} \hat{\alpha}_k$ is given by
\[
\hat{\alpha}_0=\mathcal{H}_p\alpha_0,~\hat{\alpha}_k=d^*_p G_p\sum_{i+j=k} \mathcal{L}_{\phi_i}^{1,0}\hat{\alpha}_{j}\in F^{p}A^{p+q}(X),\quad k>0.
\]
We say \textit{the deformation of filtered $(p,q)$-forms on $X$ are (canonically) unobstructed} (w.r.t. $\pi$) if for any $\alpha_0\in\ker d\cap F^{p}A^{p+q}(X)$
there is a (canonical) deformation of $\alpha_0$ on $B$.

At last, a \emph{deformation $\alpha(t)$ of $\alpha_0$} on $T$ (w.r.t. $\pi$) is said to be \emph{faithful} if $\alpha(t)$ satisfies the following: $\alpha_0\in\im d\cap F^{p}A^{p+q}(X)$ whenever $\alpha(t)\in F^pA^{p+q}(X)\cap\im d_{\phi(t)}$ for some $t\in T$.
\end{definition}
\begin{corollary}\label{Coro-123}
The following holds:
\begin{enumerate}
  \item $\dim F^pH^{k}(X_t)$ is independent of $t\in B$ if and only if the deformations of filtered $(p,k-p)$-forms and filtered $(p,k-1-p)$-forms are canonically unobstructed;
  \item For any $0\leq p\leq k$, the alternating sum $\sum_{i=0}^k \dim (-1)^{k-i}F^pH^{i}(X_t)$ and $\dim E^{0,q}_1(X_t)/E^{0,q}_\infty(X_t)$ are upper semi-continuous function of $t\in B$ (in analytic Zariski topology). Moreover, the deformations of filtered $(p,k-p)$-forms are canonically unobstructed if and only if $\sum_{i=0}^k \dim (-1)^{k-i}F^pH^{i}(X_t)$ is independent of t;
  \item The notion of canonically unobstructedness for filtered forms are independent of the choices of Hermitian metrics.
\end{enumerate}
\end{corollary}
\begin{proof}$(1)$ By the definition of canonically unobstructedness, the deformations of filtered $(p,k-p)$-forms are canonically unobstructed iff
\[
V_{t}^{p,k}=\ker\triangle_{p}\cap F^{p}A^{k}(X),\quad\text{for~any}~t\in B,
\]
which is equivalent to $v^{p,k}_t=0$ for any $t\in B$. The conclusion then follows from \eqref{eq-jumping-formula}.

$(2)$ By considering the alternating sum of \eqref{eq-jumping-formula}, we have
\[
v^{p,k}_t=\sum_{i=0}^k \dim (-1)^{k-i}F^pH^{i}(X)-\sum_{i=0}^k \dim (-1)^{k-i}F^pH^{i}(X_t),
\]
which combined with \eqref{eq-Vt-vt} implies
\[
\sum_{i=0}^k \dim (-1)^{k-i}F^pH^{i}(X_t)=\dim V_{t}^{p,k}.
\]
The upper semi-continuity of $\sum_{i=0}^k \dim (-1)^{k-i}F^pH^{i}(X_t)$ follows from this and the fact that $\dim V_{t}^{p,k}$ is upper semi-continuous. As a consequence, the upper semi-continuity of $\dim E^{0,q}_1(X_t)/E^{0,q}_\infty(X_t)$ follows from \eqref{eq-E0q}.

$(3)$ follows directly from $(2)$ since $\dim F^pH^{i}(X_t)$ are complex structure invariants.
\end{proof}

\begin{proposition}\label{prop-faithful}
Assume the deformations of filtered $(p,k-1-p)$-forms are canonically unobstructed. Let $\alpha(t)$ be a canonical deformation of $\alpha_0\in\ker d\cap F^{p}A^{k}(X)$ on $T$ (w.r.t. $\pi$), then it is faithful.
\end{proposition}
\begin{proof}
We may assume $\alpha(t)$ is given by \eqref{eq-canonical}. If $\alpha(t)\in F^pA^{p+q}(X)\cap\im d_{\phi(t)}$ for some $t\in T$, then we have $\sum_{k\geq0} \hat{\alpha}_k\in \im d_{\phi(t)}\cap\ker d_p^*$. But since the deformations of filtered $(p,k-1-p)$-forms are canonically unobstructed, we have $v^{p,k-1}_t=0$ for any $t\in B$ which implies
\[
\im d_{\phi(t)}\cap\ker d_p^*=0,\quad\text{for~any}~t\in B.
\]
Hence, we get $\sum_{k\geq0} \hat{\alpha}_k=0\Rightarrow\hat{\alpha}_0=\mathcal{H}_p\alpha_0$ by using $(3)$ of Proposition \ref{prop-unique-power-seris-solution}.
\end{proof}

\section{Deformations of Hodge filtrations}

\subsection{Faithful deformations of filtered forms}

The proof of \cite[Thm.\,3.4]{WX23} essentially implies the following:
\begin{theorem}\label{thm-WX23} Let $\pi: (\mathcal{X}, X)\to (B,0)$ be a complex analytic family over a small polydisc $B\subset\C^m$ such that for each $t\in B$ the complex structure on $X_t$ is represented by a Beltrami differential $\phi(t)$. Assume
  \[
  F^{p}A^{p+q+1}(X)\cap dA^{p+q}(X) =dF^{p}A^{p+q}(X).
  \]
Then the deformations of filtered $(p,q)$-forms are canonically unobstructed.
\end{theorem}

\begin{theorem}\label{thm-exp-FpH}
Let $\pi: (\mathcal{X}, X)\to (B,0)$ be a complex analytic family over a small polydisc $B\subset\C^m$ such that for each $t\in B$ the complex structure on $X_t$ is represented by a Beltrami differential $\phi(t)$. Assume
  \[
  F^{p}A^{p+q+1}(X)\cap dA^{p+q}(X) =dF^{p}A^{p+q}(X),
  \]
and
  \[
  F^{p}A^{p+q}(X)\cap dA^{p+q-1}(X) =dF^{p}A^{p+q-1}(X).
  \]
For any $t\in B$ and $p\leq k$, the exponential operator induces the following isomorphism of vector spaces (still denoted by $e^{i_{\phi(t)}}$)
\begin{equation}\label{eq-exponential-coho}
e^{i_{\phi(t)}}: F^pH^{p+q}(X)\longrightarrow F^pH^{p+q}(X_t):[\alpha_0]\longmapsto [e^{i_{\phi(t)}}\alpha(t)],
\end{equation}
where $\{\alpha(t)\}_{t\in B}$ is the canonical deformation of $\alpha_0$.
\end{theorem}
\begin{proof}First of all, note that \eqref{eq-exponential-coho} is the composition of the following two mappings
\begin{equation}\label{eq-can-def-map}
F^{p}H^{p+q}(X)\longrightarrow F^{p}H^{p+q}_{d_{\phi(t)}}(X):[\alpha_0]\longmapsto [\alpha(t)],
\end{equation}
where $\alpha(t)$ is the canonical deformation of $\alpha_0$ and
\begin{equation}\label{eq-dphi-dXt}
F^{p}H^{p+q}_{d_{\phi(t)}}(X)\longrightarrow F^{p}H^{p+q}(X_t):[\alpha]\longmapsto [e^{i_{\phi(t)}}\alpha].
\end{equation}
First, according to Theorem \ref{thm-WX23}, the deformations of filtered $(p,q)$-forms and filtered $(p,q-1)$-forms are canonically unobstructed. By Proposition \ref{prop-faithful}, all canonical deformations of filtered $(p,q)$-forms in this case are faithful. These two facts imply the homomorphism given by \eqref{eq-can-def-map} is well-defined and injective. The homomorphism given by \eqref{eq-can-def-map} is an isomorphism in view of \eqref{eq-wdc-dc}. So we see that \eqref{eq-exponential-coho} is well-defined and injective. In particular, we have
\[
\dim F^pH^{p+q}(X)\leq \dim F^pH^{p+q}(X_t),\quad\text{for~any}~t\in B.
\]
But according to Proposition \ref{prop-upper-semicontinuous}, $\dim F^pH^{\bullet}(X_t)$ is upper semi-continuous which implies
\[
\dim F^pH^{p+q}(X)= \dim F^pH^{p+q}(X_t),\quad\text{for~any}~t\in B.
\]
Therefore, \eqref{eq-exponential-coho} is an isomorphism.
\end{proof}

In the context of Theorem \ref{thm-exp-FpH}, the \emph{period map} can be defined as follows
\begin{equation}\label{eq-period-map}
\Phi^{p,k}:B\longrightarrow \G (f^{p,k}, H^{k}(X,\C)),\quad t\longmapsto F^pH^{k}(X_t),
\end{equation}
where $f^{p,k}:=\dim F^pH^{k}(X)$. The period map $\Phi^{p,k}$ has the following properties (c.f.~\cite[Chp.\,10]{Voi02I} for the case of K\"ahler manifolds),

\begin{theorem}\label{thm-period-map}
Assume
  \[
  F^{p}A^{p+q+1}(X)\cap dA^{p+q}(X) =dF^{p}A^{p+q}(X),
  \]
and
  \[
  F^{p}A^{p+q}(X)\cap dA^{p+q-1}(X) =dF^{p}A^{p+q-1}(X).
  \]
The following holds:
\begin{enumerate}
  \item The period map $\Phi^{p,k}$ is holomorphic;
  \item Griffiths transversality: the tangent map
  \[
  d\Phi^{p,k}_0:T_0B\longrightarrow  \Hom(F^pH^{k}(X), H^{k}(X,\C)/F^pH^{k}(X))
  \]
  has values in $\Hom(F^pH^{k}(X), F^{p-1}H^{k}(X)/F^pH^{k}(X))$;
\end{enumerate}
If furthermore $\mathcal{H}_p=\mathcal{H}_{p-1}\mid_{F^{p}A^{\bullet}(X)}$ and $\dim B=1$, then the tangent map of $\Phi^{p,k}$ is given by
\[
d\Phi^{p,k}_0(\frac{\p}{\p t}\Big|_{t=0})([\alpha_0])=[\mathcal{H}_{p-1}i_{\phi_1}\alpha_0]\in F^{p-1}H^{k}(X),\quad [\alpha_0]\in F^pH^{k}(X),
\]
where $t$ is a local coordinate function on $B$ and $\phi_1$ is the first order coefficient in $\phi(t)=\sum_{j\geq 1}\phi_jt^j$.
\end{theorem}
\begin{proof}
These statements can be deduced from Theorem \ref{thm-exp-FpH}. In fact, for any $t\in B$ we now have
\[
\Phi^{p,k}(t)=F^pH^{k}(X_t)=\C\Big\{ [e^{i_{\phi(t)}}\alpha^l(t)] \Big\}_{l=1}^{f^{p,k}},
\]
where $\{[\alpha^l_0]: l=1,\cdots,f^{p,k}\}$ is a basis of $F^pH^{k}(X)$
and $\alpha^l(t)$ is the canonical deformation of $\alpha^l_0$. This shows $\Phi^{p,k}$ is holomorphic because both $\phi(t)$ and each $\alpha^l(t)$ are holomorphic. This is $(1)$.

Next, let $\alpha(t)$ be the canonical deformation of some $[\alpha_0]\in F^pH^{k}(X)$ such that $\alpha_0\in\ker\triangle_{p}\cap F^{p}A^{k}(X)$, i.e. $\alpha_0$ is a harmonic filtered $(p,k-p)$-form. Without loss of generality assume $\dim B=1$ and we can write $\phi(t)=\sum_{j\geq 1}\phi_jt^j$, $\alpha(t)=\sum_{k\geq 0}\alpha_kt^k$, where $\alpha_{k}=d^*_p G_p\sum_{i+j=k}\mathcal{L}_{\phi_i}^{1,0}\alpha_{j}$. So we have
\[
\frac{\p}{\p t}\Big|_{t=0}\alpha(t)=\alpha_1=d^*_p G_p\mathcal{L}_{\phi_1}^{1,0}\alpha_{0}~.
\]
Now if we differentiate $e^{i_{\phi(t)}}\alpha(t)$ with respect to $t$ at $t=0$, we get
\begin{equation}\label{eq-differentiate}
\frac{\p}{\p t}\Big|_{t=0}[e^{i_{\phi(t)}}\alpha(t)]=[i_{\phi_1}\alpha_0+d^*_p G_p\mathcal{L}_{\phi_1}^{1,0}\alpha_{0}]\in F^{p-1}H^{k}(X).
\end{equation}
This is $(2)$.

Finally, if $\dim B=1$ then we can write $\phi(t)=\sum_{j\geq 1}\phi_jt^j$ with $t\in B\subset\C$ and
\[
\kappa(\frac{\p}{\p t}\Big|_{t=0})=\frac{\p}{\p t}\Big|_{t=0}\phi(t)=\phi_1.
\]
Let $[\alpha_0]\in F^pH^{k}(X)$ and $\alpha(t)=\sum_{k\geq 0}\alpha_kt^k$ a canonical deformation of $\alpha_0\in F^pA^{k}(X)\cap\ker d$. From $(1),(2)$, we know that locally,
\begin{align*}
d\Phi^{p,k}_0(\frac{\p}{\p t}\Big|_{t=0})([\alpha_0])&=\frac{\p}{\p t}\alpha(t)\Big|_{t=0}\\
&=[i_{\phi_1}\alpha_0+d^*_p G_p\mathcal{L}_{\phi_1}^{1,0}\alpha_{0}]\\
&=[\mathcal{H}_{p-1}i_{\phi_1}\alpha_0+\mathcal{H}_{p-1}d^*_p G_p\mathcal{L}_{\phi_1}^{1,0}\alpha_{0}]\\
&=[\mathcal{H}_{p-1}i_{\phi_1}\alpha_0+\mathcal{H}_{p}d^*_p G_p\mathcal{L}_{\phi_1}^{1,0}\alpha_{0}]\\
&=[\mathcal{H}_{p-1}i_{\phi_1}\alpha_0]\in F^{p-1}H^{k}(X).
\end{align*}
\end{proof}

Note that $\{e^{i_{\phi(t)}}\alpha^l(t)\}_{l=1}^{f^{p,k}}$ is a holomorphic frame of the Hodge bundle $\mathcal{F}^p:=\bigcup_{t\in B}F^pH^{k}(X_t)$ (c.f.\cite{LS18}). On the other hand, it is well-known that for K\"ahler manifolds, the following diagram is commutative (see e.g. \cite[Thm.\,10.21]{Voi02I})):
 \[
  \xymatrix{
  T_0B \ar[r]^-{\kappa}\ar[dr]_-{d\Phi^{p,k}_0} &  H^{1}(X,T_X^{1,0})  \ar[d]^-{\iota}   \\
          & \Hom(F^pH^{k}(X), F^{p-1}H^{k}(X)/F^pH^{k}(X)) , }
 \]
where $\kappa$ is the Kodaira-Spencer map and $\iota$ is defined as follows: for any $[\varphi]\in H^{1}(X,T_X^{1,0})$, $\iota([\varphi])=i_\varphi$. In our case, the homomorphism $\iota$ may be not well-defined.

\bibliographystyle{alpha}
\bibliography{reference}
\end{document}